the population of silver city was 347 when silver was discovered.ten years later the population was 6780 assuming exponential growth what was the population five years after the discovery of silver (y=ae^kx)
Answer by stanbon(75887)

the population of silver city was 347 when silver was discovered.ten years later the population was 6780 assuming exponential growth what was the population five years after the discovery of silver (y=ae^kx)
-----------
y = ab^x
---
347 = ab^0
6780 = ab^10
-------------------
From the 1st equation a = 347
Divide and solve for "b":
19.54 = b^10
b = 19.54^(1/10)
----
Equation:
f(x) = 347*10.54^(x/10)
----
What was the population five years after the discovery of silver?
f(5) = 347*19.54^(5/10)
f(5) = 347*sqrt(19.54) = 1534
=================================
cheers,
Stan H.